MAT 1322 3X : Calcul II

Summer 2007

Teacher:  Benoît Valiron

585 King Edward, room 204 Telephone: 562-5800 ext. 2104 Email: bvali087@uottawa.ca

Textbook:  Calculus, Single and Multivariable, by Hughes-Hallett and al.,

4th Edition, publisher Wiley and Sons.

Website:  http://aix1.uottawa.ca/∼bvali087/mat1322

You are encouraged to check the website regularly since news and handouts will be posted here.

Lectures:  Tuesday and Thursday, 7pm–9pm (CBY D207)
DGD:  Tuesday or Thursday, 5:30pm–7pm (CBY D207)

The two DGD sessions are identical. You may pick the one which is more convenient to you.

Office hours: Tuesday, 4pm–5:30pm, or by appointment, in KED B07B

Note: The room is not my office.

Evaluation: There will be homeworks, two partial exams in class and a 3-hours closed-book final examination. Your grade for the course is computed as follows: each test is worth 15%, the homeworks are worth 15% in total and the final exam is worth 55%. Please note:

• If your mark at the final examination is less than 40% or more than 90%, then your grade for the course is equal to the mark for the final examination.

• If you miss a test for a valid reason (approved by the teacher), its weight will be transfered to the final exam.

Help Center: A help center is available to you in case of unresolved issues with the material. From May 8 to August 20, the help center will be open from 10am to 3pm Monday to Thursday in Marion 021. The center will open the following Fridays: July 6, July 13, August 10, August 17.

Note: You are strongly advised to attend all lectures and to take notes. The material might be presented in a way that is different from the textbook. Also, the order might change. The DGD’s are there to help you understand the material and the proposed exercises are similar of what you will be asked to solve in homeworks and tests.

Important dates:

May 1  First day of class.  July 5 Test #2 (In class).
May 31  Test #1 (In class).  July 12 Last class.
June 11-15 Study break – No classes. July 19 7pm–10pm: Final exam, in MNT 202.